OpenStudy (anonymous):

8/9 times three and 3/4

OpenStudy (anonymous):

\[3\frac{3}{4}\] Can you convert that to an improper fraction? You need to do this before multiplying.

OpenStudy (anonymous):

yeah .. that is ryt

OpenStudy (anonymous):

there's a 3 in front of the 3/4

OpenStudy (anonymous):

\[\frac{8}{9}*3\frac{3}{4}\]

OpenStudy (anonymous):